OpenStudy (anonymous):

To solve by completing the square, what value should be added to both sides of the equation x^2+4x=-1??????

OpenStudy (amistre64):

how would you complete this square? \begin{array}c x^2&x&x\\ x&.&.\\ x&.&.\\ \end{array} in other words, how many places would you have to fill in?

OpenStudy (anonymous):

ohhh, i see...sortof. so just four, right? but where would the number go around the box?

OpenStudy (amistre64):

the number is just the number of places to make it a complete square; and yes its 4 :) \begin{array}c x^2&x&x\\ x&1&2\\ x&3&4\\ \end{array} the algebra of it is this; we take the (4x) part and split it in 2, so that we can make a frame work to fill in the square. this gives us a 2x2 are to fill; 2*2 = 4 spose we had x^2 + 6x; we would apply the same concepts; split the (6x) in half and construct a frame work to fill in; 3x up top and 3x down the side; we have a 3x3 area to fill in so... 3*3 = 9 for that one

OpenStudy (anonymous):

nevermind figured it out =D

OpenStudy (anonymous):

THANKS THOUGH =D

OpenStudy (amistre64):

youre welcome :)

OpenStudy (amistre64):

spose we had something a little trickier to work with like x^2 + 5x; we can still find a number to complete the square, but the picture begins to get too complicated to be useful, but the algebra still applies take 5 and split it in half: 5/2 now we need to fill in a (5/2)x(5/2) area so: (5/2)*(5/2) = 25/4 as a result